Which ordered pair is not a solution of f ( x ) = 2 x + 9?
(0, 9)
(7, 23)
(3, 14)
(2, 13)

Answer by ReadingBoosters(3246)

Using
f(x) = 2x + 9
...
Substitute each x value from the pair and confirm which y is not given.
...
2(0) + 9 = 9: (0,9) Good
2(7) + 9 = 14+9 = 23: (7,23) Good
2(3) + 9 = 6+9=15: (3,15) Oops, (3,14) is not an ordered pair!
2(2) + 9 = 4+9=13: (2,13) Good
.....................
Delighted to help.
-Reading Boosters
Wanting for others what we want for ourselves.
www.MyHomeworkAnswers.com